# Cluster Categorification of Rank 2 Webs

### v1

@@ Acknowledgements
This project started during the Junior Trimester Programme: New Trends in Representation Theory at the Hausdorff Institute for Mathematics in Bonn, 2020. We thank Gustavo Jasso and Jan Schröer for organising the programme, and the Hausdorff Institute for their support. Parts of this work is completed at the Cluster algebras and representation theory programme at the Isaac Newton Institute for Mathematical Sciences in 2021 (supported by EPSRC grant no EP/R014604/1). EY is supported in part by the Royal Society Wolfson Award RSWF/R1/180004.
2. Webs of Rank $2$
@@
A $k$-web is a bipartite graph drawn in a disc with weights on the edges such that the sum of weights around each interior vertex is $k$. We will be interested in $k$-webs that are inside a disc with $n$ vertices on the boundary, all colored black, numbered $1$ to $n$. Denote $[n]:=\{1,2,\cdots,n\}.$
@@
Any $k$-web will have some number edges ending at the boundary vertices. Let us call such edges leaves. Note that leaves can intersect each other inside the disk or on the boundary vertices. We will restrict our attention to webs such that the leaves all have weight $1$. A boundary vertex can have multiple (or no) leaves attached to it. Any such $k$-web gives a well-defined function on $\rm Gr(k,n)$ [FLL19].
@@
For any $k$-web with with all leaves of weight $1$, the number of leaves is a multiple of $k$. We will call a $k$-web with $rk$ leaves a $k$-web (or simply a web) of rank $r$.
@@
In this paper, we will only consider rank $1$ and rank $2$ webs. Rank $1$ webs consist of one interior white vertex and $k$ boundary black vertices. The rank $2$ webs $W$ we consider will consist of a tree with one black vertex and three white vertices inside the disk and edges as shown below: White vertices are connected to the boundary vertices with finite number of edges.
Figure 1. An illustration of rank $1$ and $2$ webs.
@@
The set $[n]$ has a natural cyclic structure. For any $i\in[n]$ the cyclic order induces a total order $<_{i}$, where $i<_{i}i+1<_{i}\cdots<_{i}n<_{i}1<_{i}\cdots<_{i}i-1$. Let $R,S,T$ be three subsets of $[n]$. We will say that they are cyclically separated if there is some $i$ such that all the elements of $R$ are less than all the elements of $S$, which are in turn less than all the elements of $T$ in the ordering $<_{i}$. In particular, the cyclically separated sets are disjoint.
@@
Let $R,S,T$ be cyclically separated subsets of $[n]$ and let $V\subset[n]$ be an arbitrary (possibly empty) set disjoint from $R,S$ and $T$. Suppose further that $|R|=a,|S|=b,|T|=c,|V|=d$ and $a+b+c+2d=2k$. Then we can define a rank $2$ web $W(R,S,T,V)$ as follows: The elements in $R,S$, $T$ and $V$ will be the labels of boundary vertices that are leaves of the tree. The vertices in $R$ will be attached to the first white vertex; the vertices in $S$ will be attached to the second white vertex, and the vertices of $T$ will be attached to the third white vertex. Each vertex of $V$ will be attached to two white vertices, but which white vertices they are attached to inside the disk does not matter. Changing the attachments of the vertices in $V$ will give us equivalent webs, in the sense that they give the same function on the Grassmannian, see [FP16].
@@
Consider the homogeneous coordinate ring $\mathbb{C}[\rm Gr(3,6)]$. The cluster algebra structure found in this ring is of Dynkin type $D_{4}$; twenty-two cluster variables of which we have six frozen variables. Since all Plücker coordinates are cluster variables, we have $\binom{6}{3}=30$ of them. This means we have two non Plücker cluster variables in this case. They can be found using exchange relations, or skein relations on webs [FP16].
@@
The set $\mathcal{W}$ is the complete set of rank $2$ cluster variables on the Grassmannian $\rm Gr(k,n)$.
Strong evidence for this is the work of Fraser [Fra22] which describes the full canonical basis of rank $2$ functions as webs. Among those webs, the only arborizable ones are those in $\mathcal{W}$. It is believed that a web must be arborizable in order to be a cluster variable [FP16]. (They conjecture this for $k=3$, but we believe it has a good chance of being true for $k>3$ as well.) Note that the number of webs in $\mathcal{W}$ is also precisely the cardinality of rank $2$ indecomposable modules predicted in [BBL21].
@@ 3. JKS Modules
We follow the notation of Jensen-King-Su [JKS16]. Their paper gives an additive categorification of the cluster algebra structure in $\mathbb{C}[\rm Gr(k,n)]$ by using certain Cohen-Macaulay (CM) modules. We will be interested in rank $1$ and rank $2$ CM modules defined in [JKS16]. Let us give a gentle introduction to this categorification. Fix $k,n.$ Let $Q$ be a quiver as shown in Figure 3. Let $\mathcal{A}_{k,n}$ be the path algebra of $Q$ with the relations ‘$y_{i+1}x_{i+1}=x_{i}y_{i}$’ and ‘$x_{i+k}\cdots x_{i+2}x_{i+1}=y_{i-(n-k-1)}\cdots y_{i-1}y_{i}$’ for each vertex $i$. We consider $1,\cdots,n$ cyclically, so $0=n$ and so on. Notice that we have $2n$ relations in total.
@@
For the simplicity of the presentation, we mostly draw $Q$ in a line with $n+1$ vertices thinking that vertices 00 and $n$ identified (see Figure 4).
In [JKS16], they showed that rank $1$ modules are in bijection with the $k$-subsets of $[n]$ which are Plücker coordinates in the corresponding Grassmannians. We use a profile for rank $1$ modules which is defined as below.
@@
A profile of a rank $1$ module is a series of up and down moves written respecting $k$-subsets of $[n]$ as down steps and the rest of $n-k$ as up steps.
@@
In Figure 4, we consider a rank $1$ module $I=\{1,2,5\}$ for $Gr(3,6)$. We describe the contours of rank $1$ modules as a sequence of $n-k$ $U$’s (up) and $k$ $D^{\prime}s$ (down). The contour of $I=\{1,2,5\}$ in $Gr(3,6)$ is $DDUUDU.$
@@
Note that there are regions appearing between the contours in a rank $2$ module. We will call these regions “boxes” (a similar definition was given in [BBGEL22, Definition 4.1]).
@@
Let $\frac{I}{J}$ be a profile consisting of three boxes with $(\frac{U}{U})$ and $(\frac{D}{D})$ interspersed. Then there is a unique indecomposable module of rank $2$ with this profile. Such rank $2$ modules are the modules in $\mathcal{M}$.
@@
It is conjectured in [BBGE20, Conjecture 5.15] that there are $2\binom{n}{6}\binom{n-6}{k-3}$ rigid indecomposable rank $2$ modules corresponding to real roots. In the same paper it is shown that ${\rm CM}(B_{k,n})$ has at most $2\binom{n}{6}\binom{n-6}{k-3}$ rigid indecomposable rank $2$ modules which correspond to real roots. In [BBGEL22, Theorem 4.7], they show that there are at least
(1) $\leavevmode\nobreak\ N_{k,n}:=\sum^{k}_{r=3}(\frac{2r}{3}p_{1}(r)+2rp_{2}(r)+4rp_{3}(r))\binom{n}{2r}\binom{n-2r}{k-r}$
rigid rank $2$ indecomposable modules that correspond to roots and conjecture that this is all of them. This is precisely the number that we construct.
We also know from [BBGE20, Proposition 5.7] that if $M$ is a rank $2$ module with profile $\frac{I}{J}$ and $M$ corresponds to a real root, then $|I\cap J|=k-3$. By stretching the modules corresponding to the web in Figure 2, we obtain $2\binom{n}{6}\binom{n-6}{k-3}$ modules corresponding to real roots, verifying [BBGE20, Conjecture 5.15].
@@ 4. Webs and Modules
In a rank $2$ web, we have three white vertices inside the disk and each white vertex connected to several vertices on the boundary. As before, there will be $a$ vertices on the boundary will connect to one white vertex, $b$ vertices will connect to the second white vertex, and $c$ vertices will connect with the third white vertex. There will be $d$ vertices that are attached to any two white vertex. Set $a=\alpha+\beta$, $b=\beta+\gamma$ and $c=\alpha+\gamma$ and solve for $\alpha,\beta$ and $\gamma$. Set
@@
Assume $d=0$, then we can visualize this map as follows (update the fig).
@@
Note that the count of the modules corresponding to real roots is $2\binom{12}{6}\binom{6}{2},$ corresponding to the first term in the the sum.
@@
We may now state the main theorem of the paper:
@@
If $W\in\mathcal{W}$, then $W$ is a cluster variable, and the corresponding module is $\Psi(W)\in\mathcal{M}.$
@@ 5. Stretching Functions
We can define these functions on modules of any rank by just applying them to each contour in the profile of the module. For example, $U_{i}(\frac{I}{J})=\frac{U_{i}(I)}{U_{i}(J)}$.
We have described the effect of $U_{i}$ and $D_{i}$ on the contours of a module. It is simple to see how to extend $U_{i}$ and $D_{i}$ to functors on categories of modules: Consider a module $M$. Then $U_{i}(M)$ is the module with $(U_{i}(M))_{j}=M_{j}$ for $j\leq i$ and $(U_{i}(M))_{j}=M_{j-1}$ for $j>i$. The maps $x$ and $y$ are defined as before, except that $y_{i}$ is the identity map on $M_{i}$, and $x_{i}$ is the map given by multiplication by $t$. This is determined by the relation $y_{i+1}x_{i+1}=t=x_{i}y_{i}.$
@@
Here is an example of stretching of a rank $2$ web and a rank $2$ module under the up strectching map $U_{3}$.
@@
The following lemma shows that modules corresponding to cluster variables behave while under stretching.
@@
Suppose that $M$ is an indecomposable and rigid module corresponding to a cluster variable $f$. Then $U_{i}(M)$ and $D_{i}(M)$ for each $i\in[n]$ are indecomposable and rigid modules corresponding to the functions $U_{i}^{*}(f)$ and $D_{i}^{*}(f)$, respectively.
@@
For any cluster variable $f\in C$, $F_{i}^{*}(f)$ will be a cluster variable on $Gr(k,n+1)$. The set $\{F_{i}^{*}(f)|f\in C\}$ can be completed to a cluster for $Gr(k,n+1)$, call this cluster $C^{\prime}$. Then the quiver for the cluster $C$ is a subquiver of the quiver for $C^{\prime}$. For $f\in C$, we know that if $M$ is the module corresponding to $f$, then $U_{i}(M)$ is the module corresponding to $F_{i}^{*}(f)$. This is because $f$ was chosen to be a Plücker coordinate.
Then we need to use the following lemma:
@@
Suppose that in a cluster $C$ on $Gr(k,n)$. If $f\in C$ corresponds to a module $M$, then $F_{i}^{*}(f)$ corresponds to the module $U_{i}(M)$. Let $f^{\prime}$ be the mutation of $f$, having corresponding module $M^{\prime}$. Then $F_{i}^{*}(f^{\prime})$ corresponds to the module $U_{i}(M^{\prime})$.
This comes down to the fact that if we have an exact sequence
$M^{\prime}\rightarrow M^{\prime\prime}\rightarrow M$
of $\mathcal{A}_{k,n}$-modules, then
$U_{i}(M^{\prime})\rightarrow U_{i}(M^{\prime\prime})\rightarrow U_{i}(M)$
is an exact sequence of of $\mathcal{A}_{k,n+1}$-modules. This is similarly true if $M$ and $M^{\prime}$ are switched. This suffices to prove the lemma.
We now return to the proof of Lemma 5.4. We have seen that if the compatibility between $U_{i}$ and $F_{i}^{*}$ holds for all the functions in the cluster $C$, it holds for all the functions in any mutation of $C$. Hence, by recursion, it holds for all cluster variables. ∎

### v2

@@ Acknowledgements
We thank Alastair King for his valuable comments on the first draft of this paper and pointing out the use of immanants (Remark 3.3). This project started during the Junior Trimester Programme: New Trends in Representation Theory at the Hausdorff Institute for Mathematics in Bonn, 2020. Part of this work is completed at the CAR programme at the Isaac Newton Institute for Mathematical Sciences in 2021 (supported by EPSRC grant no EP/R014604/1). We thank organisers and both institute for their support. EY is supported by different bodies: in part by the Royal Society Wolfson Award RSWF/R1/180004 and by the Ministry of Education and Science of the Republic of Bulgaria with grant no DO1-239/10.12.2024 and the Simons Foundation with grant no SFI-MPS-T-Institutes-00007697.
2. Webs of Rank 22
@@
A $k$-web is a bipartite graph drawn in a disc with weights on the edges such that the sum of weights around each interior vertex is $k$. We will be interested in $k$-webs that are inside a disc with $n$ vertices on the boundary, all colored black, numbered $1$ to $n$. Denote $n]:=\{1,2,\cdots,n\}.$
@@
Any $k$-web will have some number of edges ending at the boundary vertices. Let us call such edges leaves. Note that leaves can intersect each other inside the disk or on the boundary vertices. We will restrict our attention to webs such that the leaves all have weight $1$. A boundary vertex can have multiple (or no) leaves attached to it.
@@
For any $k$-web with all leaves of weight $1$, the number of leaves is a multiple of $k$. We will call a $k$-web with $rk$ leaves a $k$-web (or simply a web) of rank $r$.
@@
In this paper, we will only consider rank $1$ and rank $2$ webs. Rank $1$ webs consist of one interior white vertex and $k$ boundary black vertices. The rank $2$ webs $W$ we consider will consist of a tree with one interior black vertex and three white vertices inside the disk and edges as shown below: White vertices are connected to the boundary vertices with a finite number of edges.
Figure 1. An illustration of rank $1$ and 22 webs.
@@
The set $[n]$ has a natural cyclic structure when placed on a boundary of a circle. For any $i\in[n]$ the cyclic order induces a total order $_{i}$, where $i<_{i}i+1<_{i}\cdots<_{i}n<_{i}1<_{i}\cdots<_{i}i-1$. Let $R,S,T$ be three subsets of $[n]$. We will say that they are cyclically separated if there is some $i$ such that all the elements of $R$ are less than all the elements of $S$, which are in turn less than all the elements of $T$ in the ordering $_{i}$. In particular, the cyclically separated sets are disjoint.
@@
Let $R,S,T$ be cyclically separated subsets of $[n]$ and let $V\subset[n]$ be an arbitrary (possibly empty) set disjoint from $R,S$ and $T$. Suppose further that $|R|=a,|S|=b,|T|=c,|V|=d$ and $a+b+c+2d=2k$. Then we can define a rank $2$ web $W(R,S,T,V)$ as follows: The elements in $R,S$, $T$ and $V$ will be the labels of boundary vertices that are leaves of the tree. The vertices in $R$ will be attached to the first white vertex; the vertices in $S$ will be attached to the second white vertex, and the vertices of $T$ will be attached to the third white vertex. Each vertex of $V$ will be attached to two white vertices, but which white vertices they are attached to inside the disk does not matter. Changing the attachments of the vertices in $V$ will give us equivalent webs, in the sense that they give the same function on the Grassmannian, see [FP16, Section 5].
@@
A $k$-web of rank $r$ on the vertex set $[n]$ induces a degree $r$ function on the homogeneous coordinate ring, $\mathbb{C}[\rm Gr(k,n)]$, of the Grassmannian $\rm Gr(k,n)$: Such a web gives a function of $n$ vectors in a $k$-dimensional vector space; moreover, this function is invariant under the action of $SL_{k}$. Therefore, these invariant functions in $\mathbb{C}[\rm Gr(k,n)]$ can be understood by viewing a point in the Grassmannian as a $k\times n$ matrix and evaluating the invariant on the columns of the matrix.
For example, recall that a Plücker coordinate in $\mathbb{C}[\rm Gr(k,n)]$ is an ${\rm SL}_{k}$- invariant function of $k$ vectors. This is represented by $k$-valent graph shown in Figure 1 (left). Throughout this paper, we will therefore view webs as functions on the Grassmannian. For a detailed treatment of the subject, we refer the reader to [FLL19, Example 3.4], [FP16, Section 5].
@@
Consider the homogeneous coordinate ring $\mathbb{C}[\rm Gr(3,6)]$. The cluster algebra structure found in this ring is of Dynkin type $D_{4}$; twenty-two cluster variables of which we have six frozen variables. Since all Plücker coordinates are cluster variables, we have $\binom{6}{3}=20$ of them. This means we have two non-Plücker cluster variables in this case. They can be found using exchange relations, or skein relations on webs [FP16].
@@
The set $\mathcal{W}$ realizes the complete set of rank $2$ cluster variables on the Grassmannian $\rm Gr(k,n)$.
Strong evidence for this is the work of Fraser [FRA22] which describes the full canonical basis of rank $2$ functions as webs. Among those webs, the only arborizable ones are those in $\mathcal{W}$. Informally speaking, arborization is a process that turns a planar web into a possibly self-intersecting tree diagram if possible, see [FP16, Definition 10.3]. It is believed that a web must be arborizable in order to be a cluster variable [FP16]. (They conjecture this for $k=3$, but we believe it has a good chance of being true for $k>3$ as well.) Note that the number of webs in $\mathcal{W}$ is also precisely the number of isomorphism classes of rank $2$ indecomposable rigid modules predicted in [BBL21].
@@ 3. JKS Modules
We follow the notation of Jensen-King-Su [JKS16]. Their paper gives an additive categorification of the cluster algebra structure in $\mathbb{C}[\rm Gr(k,n)]$ by using certain Cohen-Macaulay (CM) modules. We will be interested in rank $1$ and rank $2$ CM modules defined in [JKS16]. Let us give a gentle introduction to this categorification. Fix $k,n.$ Let $Q=(Q_{0},Q_{1})$ be a quiver with a vertex set $Q_{0}$ and an edge set $Q_{1}$ as shown in Figure 3. Let $\mathcal{A}_{k,n}$ be the path algebra of $Q$ with the relations $y_{i+1}x_{i+1}=x_{i}y_{i}$’ and $x_{i+k}\cdots x_{i+2}x_{i+1}=y_{i-(n-k-1)}\cdots y_{i-1}y_{i}$’ for each vertex $i$. We consider $1,\cdots,n$ cyclically, so $0=n$ and so on. Notice that we have $2n$ relations in total.
@@
For the simplicity of the presentation, we mostly draw $Q$ in a line with $n+1$ vertices thinking that vertices 0 and $n$ identified (see Figure 4).
In [JKS16], they showed that rank $1$ modules are in bijection with the $k$-subsets of $[n]$ which are Plücker coordinates in the corresponding Grassmannians. We use a profile, defined in [JKS16], for rank $1$ modules as below.
@@
A contour of a rank $1$ module is a series of up and down moves written respecting $k$-subsets of $[n]$ as down steps and the rest of nkn-k as up steps.
A profile of a module is a collection of contours.
@@
In Figure 4, we consider a rank $1$ module $I=\{1,2,5\}$ for $Gr(3,6)$. We describe the contours of rank $1$ modules as a sequence of nkn-k $U$’s (up) and $k$ $D^{\prime}s$ (down). The contour of $I=\{1,2,5\}$ in $Gr(3,6)$ is $DDUUDU.$
@@
Note that there are regions appearing between the contours in a rank $2$ module. We will call these regions “boxes” (a similar definition was given in [BBG+22, Definition 4.1]).
@@
Let $\frac{I}{J}$ be a profile consisting of three boxes with $(\frac{U}{U})$ and $(\frac{D}{D})$ interspersed. Then there is a unique indecomposable module of rank $2$ with this profile. All the modules in $\mathcal{M}$ are of this type.
@@
A module $M$ is called rigid if and only if $\operatorname{Ext}^{1}(M,M)=0$. It is conjectured in [BBG20, Conjecture 5.15] that there are $2\binom{n}{6}\binom{n-6}{k-3}$ rigid indecomposable rank $2$ modules corresponding to real roots. In the same paper it is shown that ${\rm CM}(A_{k,n})$ has at most $2\binom{n}{6}\binom{n-6}{k-3}$ rigid indecomposable rank $2$ modules which correspond to real roots. In [BBG+22, Theorem 4.7], they show that there are at least
(1) $\penalty 10000\ N_{k,n}:=\sum^{k}_{r=3}(\frac{2r}{3}p_{1}(r)+2rp_{2}(r)+4rp_{3}(r))\binom{n}{2r}\binom{n-2r}{k-r}$
rigid rank $2$ indecomposable modules that correspond to roots and conjecture that this is all of them. In this formula, $p_{1}(r),p_{2}(r),p_{3}(r)$ are the number of partitions of $r$ into three parts which are equal, two of which are equal, or all of which are distinct, respectively. This is precisely the number of modules that we constructed.
We also know from [BBG20, Proposition 5.7] that if $M$ is an indecomposable rank $2$ module with profile $\frac{I}{J}$ and $M$ corresponds to a real root, then $|I\cap J|=k-3$. By stretching the modules, Definition 5.1, corresponding to the web in Figure 2, we obtain $2\binom{n}{6}\binom{n-6}{k-3}$ modules corresponding to real roots, verifying [BBG20, Conjecture 5.15]. This implies that indecomposable rank $2$ modules of $\mathcal{M}$ are rigid.
Remark 3.3.
In our notation for webs, the partition corresponds to $r=a+b+c,$ where $a=|R|$, b=|S|,c=|T|,kr=|V|b=|S|,c=|T|,k-r=|V| for a web $W(R,S,T,V).$
Note that by using immanants [FRA22], we can also write
(2) $N_{k,n}:=\sum^{k}_{r=3}2\binom{r}{3}\binom{n}{2r}\binom{n-2r}{k-r}$.
Here, $2\binom{r}{3}$ counts the ways of dividing $2r$ cyclically arranged elements into three sectors of even size, which gives via the immanant construction the webs in our set $\mathcal{W}.$ Note that these sectors have sizes $2a,2b,2c$ for a web $W(R,S,T,V)$ in the notation above.
@@ 4. Webs and Modules
In a rank $2$ web, we have three white vertices inside the disk and each white vertex is connected to several vertices on the boundary. As before, on the boundary, there will be $a$ vertices connected to the first white vertex, $b$ vertices connected to the second white vertex, and $c$ vertices connected to the third white vertex. There will be $d$ vertices that are attached to any two white vertex. Set $a=\alpha+\beta$, $b=\beta+\gamma$ and $c=\alpha+\gamma$ and solve for $\alpha,\beta$ and $\gamma$. Set
@@
Assume $d=0$, then we can visualize this map as follows.
@@
Note that the count of the modules corresponding to real roots is $2\binom{12}{6}\binom{6}{2}$, corresponding to the first term in the sum.
@@
We now state the main theorem of the paper.
@@
If $W\in\mathcal{W}$, then $W$ is a cluster variable, and the corresponding module is $\Psi(W)\in\mathcal{M}.$ In particular, the cluster character of $\Psi(W)$ is $W.$
@@ 5. Stretching Functions
We would like to define stretching functions on modules. We can first define stretching on profiles by just applying stretching to each contour in the profile. For example, $U_{i}(\frac{I}{J})=\frac{U_{i}(I)}{U_{i}(J)}$.
We have described the effect of $U_{i}$ and $D_{i}$ on the contours of a module. It is simple to see how to extend $U_{i}$ and $D_{i}$ to functors on categories of modules:
Definition 5.2.
Consider a module $M$. Then $U_{i}(M)$ is the module with $(U_{i}(M))_{j}=M_{j}$ for $j\leq i$ and $(U_{i}(M))_{j}=M_{j-1}$ for $j>i$. The maps $x$ and $y$ are defined as before, except that $y_{i}$ is the identity map on $M_{i}$, and $x_{i}$ is the map given by multiplication by $t$. This is determined by the relation $y_{i+1}x_{i+1}=t=x_{i}y_{i}.$
@@
Remark 5.3.
Stretching was also defined in [BBG+22], where $U$ and $D$ were called “increase” and “co-increase.”
Remark 5.4.
It is clear that stretching preserves indecomposability of modules. It is less obvious that it preserves rigidity and reachability. We will show later in Lemma 5.8 that stretching a reachable rigid module gives another reachable rigid module.
@@
Here is an example of stretching of a rank $2$ web and a rank $2$ module under the up stretching map $U_{3}$.
@@
We will see in Lemma 5.8 that the stretching maps $U_{i}$ (respectively $D_{i}$) are induced by pullback of functions under $F_{i}$ (respectively $G_{i}$).
@@
In fact, more is true. Let $C$ be a cluster on $Gr(k,n)$ consisting of Plücker coordinates. Then pulling back under $F_{i}^{*}$ gives cluster variables on $Gr(k,n+1)$ which can be completed to a cluster $\tilde{C}$ on $Gr(k,n+1)$. Note that the variables that are added are all frozen variables, and all the arrows in the quiver for $\tilde{C}$ only connect these frozen variables with variables that were frozen in $C$. If we start mutating the cluster $C$, there is a corresponding set of mutations on the cluster $\tilde{C}$. These mutations never touch the variables frozen in $C$, so the arrows involving the new frozen variables in $\tilde{C}$ never change. Thus every cluster on $Gr(k,n)$ can be completed to one on $Gr(k,n+1)$, and the quivers for clusters on $Gr(k,n)$ are all subquivers of the corresponding clusters on $Gr(k,n+1)$. To summarize,
Proposition 5.7.
[FL19, Section 3] The following is true for any cluster $C$ on $Gr(k,n)$:
  • (i)
    Pulling back the cluster variables in $C$ via $F_{i}^{*}$ (respectively $G_{i}^{*}$) gives cluster variables on $Gr(k,n+1)$ (respectively $Gr(k+1,n+1)$) that can be completed to a cluster $\tilde{C}$ on $Gr(k,n+1)$ (respectively $Gr(k+1,n+1)$).
  • (ii)
    $C$ is a subquiver of $\tilde{C}$.
  • (iii)
    The additional arrows in $\tilde{C}$ only connect frozen vertices in $\tilde{C}$ to arrows that were frozen in $C$. Thus the additional arrows are independent of the cluster $C$.
Note that the reachable rigid indecomposable modules are in one-to-one correspondence with the cluster variables under the cluster character map [GLS08, JKS16]. Moreover, mutation of cluster variables corresponds to mutation of modules. We give below the properties that characterize the mutation of modules. The following lemma shows that modules corresponding to cluster variables behave well under stretching.
@@
Suppose that $M$ is an indecomposable, reachable rigid module corresponding to a cluster variable $f$. Then $U_{i}(M)$ and $D_{i}(M)$ for each $i\in[n]$ are indecomposable, reachable rigid modules corresponding to the functions $F_{i}^{*}(f)$ and $G_{i}^{*}(f)$, respectively.
@@
Say that a cluster $C$ behaves well under stretching if the following hold:
  1. (1)
    If $f\in C$ corresponds to a module $M$, then $F_{i}^{*}(f)$ corresponds to the module $U_{i}(M)$
  2. (2)
    Let $f$ and $g$ in $C$ correspond to modules $M$ and $N$. Then any arrow in the quiver between the vertices of $f$ and $g$ gives a corresponding irreducible map $\alpha:M\rightarrow N$. In this case, we have that $U_{i}(\alpha):U_{i}(M)\rightarrow U_{i}(N)$ is also an irreducible map.
@@
For $f\in C$, we know that if $M$ is the module corresponding to $f$, then $U_{i}(M)$ is the module corresponding to $F_{i}^{*}(f)$. This is because $f$ was chosen to be a Plücker coordinate.
Moreover, one can check that the irreducible maps between the modules corresponding to cluster variables in $C$ give irreducible maps after stretching. Thus the initial cluster $C$ behaves well under stretching.
We wish to show that the property of behaving well under stretching is stable under mutation:
@@
Suppose $C$ is a cluster for $Gr(k,n)$ which behaves well under stretching as in Proof 5. If $f^{\prime}$ is the mutation of $f$ having corresponding module $M^{\prime}$, then $F_{i}^{*}(f^{\prime})$ corresponds to the module $U_{i}(M^{\prime})$. Moreover, the irreducible maps between modules for the mutation of $C$ stretch to give irreducible maps between modules for the mutation of $\tilde{C}$.
Let $T$ be the cluster tilting object for the cluster $C$. Let $\tilde{T}$ be the cluster tilting object for the cluster $\tilde{C}$.
Because $M^{\prime}$ comes from $M$ via mutation, we get an exact sequence
$M^{\prime}\xrightarrow{\alpha}M^{\prime\prime}\xrightarrow{\beta}M.$
This exact sequence is characterized by the map $\beta$ coming from the irreducible maps in the category ${\rm End}(T)$ which are encoded in the quiver for the cluster $C$ [JKS22]. This gives that this is an exact sequence of $\mathcal{A}_{k,n}$-modules which is moreover a non-split sequence where $\beta$ is a minimal right $\rm add$}-$(T/M)$-approximation of $M$. This exact sequence then determines $M^{\prime}$ uniquely [JKS16].
Then because $C$ behaves well under stretching,
$U_{i}(M^{\prime})\xrightarrow{U_{i}(\alpha)}U_{i}(M^{\prime\prime})\xrightarrow{U_{i}(\beta)}U_{i}(M)$
is also an exact sequence where $U_{i}(\beta)$ consists of the appropriate irreducible maps dictated by the quiver for $\tilde{C}$. Thus it is also a non-split sequence, and $U_{i}(\beta)$ is a minimal right $\rm add$$(\tilde{T}/U_{i}(M))$-approximation of $U_{i}(M).$ This uniquely characterizes the module corresponding to the mutation of $F_{i}^{*}(f)$, which then must be $U_{i}(M^{\prime})$ using the results of [JKS16, JKS22], which ultimately depend on [GLS08].
After mutation, let the cluster tilting objects be $T^{\prime}$ and $\tilde{T}^{\prime}$. Let us now show that the irreducible maps are also compatible with stretching.
The irreducible maps between the irreducible summands of $T^{\prime}$ (respectively, $\tilde{T}^{\prime}$) come from irreducible maps between irreducible summands of $T$ (respectively, $\tilde{T}$), compositions of such maps, or the map $\alpha$ (respectively, $U_{i}(\alpha)$). In all cases, we see that irreducible maps are preserved under the stretching operation.
Note that all this continues to hold if the roles of $M$ and $M^{\prime}$ are switched, either by repeating the proof or using that these categories are $2$-Calabi-Yau. This completes the proof that behaving well under stretching is stable under mutation.
We now return to the proof of Lemma 5.8. We have seen that if the compatibility between $U_{i}$ and $F_{i}^{*}$ holds for the functions in the cluster $C$, it holds for the functions in any mutation of $C$. Hence, by recursion, it holds for all cluster variables. ∎